### v1

# IP-CRR: Information Pursuit for Interpretable Classification of Chest Radiology Reports

Abstract
The development of AI-based methods for analyzing radiology reports could lead to significant advances in medical diagnosis—from improving diagnostic accuracy to enhancing efficiency and reducing workload. However, the lack of interpretability in these methods has hindered their adoption in clinical settings. In this paper, we propose an interpretable-by-design framework for classifying radiology reports. The key idea is to extract a set of most informative queries from a large set of reports and use these queries and their corresponding answers to predict a diagnosis. Thus, the explanation for a prediction is, by construction, the set of selected queries and answers. We use the Information Pursuit framework to select informative queries, the Flan-T5 model to determine if facts are present in the report, and a classifier to predict the disease. Experiments on the MIMIC-CXR dataset demonstrate the effectiveness of the proposed method, highlighting its potential to enhance trust and usability in medical AI. Code is available at: https://github.com/Glourier/IP-CRR.git.
Keywords:
Interpretable machine learning Text classification Radiology reports.
1 Introduction
When a doctor makes a decision that can potentially have life-changing consequences for a patient, the transparency of the doctor’s decision-making is invaluable. Therefore, as machine learning methods become increasingly used in clinical practice, there is a pressing need to develop safe and trustworthy models. For a given prediction task, interpretability in machine learning serves as a crucial aspect of model transparency that answers the “why’s” and “how’s” of a model’s prediction, fostering trust between the user and the machine.
Traditionally, interpretability is carried out by first training a black-box, opaque model that optimizes performance, and then producing an explanation in the form of heatmaps [30, 6], selection of features [28, 21], or more generally, attributions [33, 31]. These post-hoc explanation methods have found overwhelming success in many areas such as computer vision, finance, and even healthcare. However, these methods lack faithfulness [29]. That is, the explanation could potentially be an artifact of the explanation method rather than evidence of the model’s underlying reasoning. To address this, interpretable-by-design models were proposed to make predictions solely based on a set of interpretable factors [16]. Interpretable factors take the form of a sparse set of concepts [18] or a set of questions and answers [7] about the given test sample. For example, disease diagnosis for a patient can be conducted by a list of symptoms, which indicates the explanations for the final prediction. While interpretable-by-design methods address the issue of faithfulness, they often require making careful and meticulous designs about the model architecture, optimization, and the set of interpretable concepts or questions. This complexity poses a significant challenge, particularly in the medical domain, where the need for both accuracy and interpretability is critical for clinical decision-making.
Figure 1: Example of the interpretable-by-design method.
Take the task of Chest Radiology Report (CRR) classification as an example (Figure 1). A given report typically contains a few key insights about the patient, and the task is to predict the disease that the patient presents. For the prediction to be interpretable, it must be grounded on a set of queries about the patient or the report, such as “Is the heart contour enlarged?” or “Does the lung volume decrease?”. To be able to make a prediction, we need not only the queries, but also the answers. Importantly, there might potentially exist queries that do not have readily available answers in the report. This setting drastically differs from the existing settings that interpretable-by-design methods are applied to, where all answers are typically available and easily obtained using out-of-the-shelf models [8].
In this work, we propose an interpretable-by-design framework for CRR classification, which extends the Information Pursuit (IP) framework for explainable classification of natural images [9, 7]. Our proposed IP-CRR method involves: 1) curating a set of interpretable queries by mining them from existing radiology reports; 2) applying natural language inference to determine if a fact is present in the report to be classified; and 3) learning a classifier that makes predictions by sequentially selecting a list of queries in order of information gain. Implementing our approach requires two key innovations relative to [9, 7]: 1) modifying the IP framework to accommodate large language models (LLMs) for answering queries, and 2) handling “Unknown” answers, as illustrated in Figure 1. Empirically, our method adapts to queries with readily available answers in the report by leveraging natural language inference. Moreover, our method achieves higher F1 score against black-box language baselines such as Flan-T5-large as well as interpretable-by-design methods such as Concept Bottleneck Models. Qualitatively, the explanations provided by IP-CRR offer a potential venue for trustworthy and transparent predictions in clinical decision-making.
2 Related Work
Radiology Report Classification. Automatic disease diagnosis from radiology reports is a critical task that enables large-scale use of machine learning models in clinical settings. Natural language processing methods are widely used to extract diagnostic labels from free-text radiology reports. There are two main approaches for this task: rule-based labeling and machine learning-based labeling. Rule-based methods [5, 25, 4, 14] rely on predefined rules or regular expressions to extract diseases from the report. However, these methods struggle with the complexity and variability of clinical language. Machine learning-based methods [32, 34, 26, 13] learn the patterns of text automatically without relying on handcrafted rules. For example, CheXbert [32] is a BERT-based medical domain specialized model pretrained on rule-based annotations, and finetuned on a small expert annotated dataset. However, despite their performance, they lack interpretability, making it hard for radiologists to trust their predictions in clinical use.
Interpretable-by-design Methods. Interpretable-by-design methods aim to produce explanations that are faithful to the model. State-of-the-art methods ensure this by using an intermediate layer of interpretable factors (e.g., concepts or questions) based on which a prediction is made. For example, Concept Bottleneck Models (CBMs) first map a given input to a list of concept scores, each corresponding to the presence of a concept in the input, and then linearly map concept scores to a predicted label. CBMs have found many successful applications in domains such as image classification and clinical prediction [23, 10, 35, 17, 24]. However, CBMs are sensitive to having concepts of very high quality and their use of a linear classifier limits their accuracy. In this work, we focus on another framework known as Information Pursuit (IP) [9, 7, 12], which provides interpretability via sequentially selecting queries in order of information gain, and then making predictions based on the sequence of query-answer pairs. Therefore, IP aligns more closely with the interactive question-answering process between a doctor and a patient. The challenge, however, involves constructing a set of meaningful queries and obtaining their answers. Currently, it has only been explored for the image classification domain, where [8] leverages large language models [1] and vision-language models [19]. In this work, we aim to leverage natural language inference for text and medical domain. We discuss the framework in detail in Section 3.3.
3 Method
In this section, we describe IP-CRR, the proposed interpretable-by-design model for classifying radiology reports. Our pipeline consists of three main components: 1) generate a set of queries from a large amount of reports (Sec. 3.2), 2) design a mechanism to answer them automatically (Sec. 3.2), and 3) integrate the Information Pursuit framework (Sec. 3.3) to conduct interpretable classification.
Figure 2: Framework of IP-CRR, which consists of three parts: 1) query set generation by extracting and clustering facts from a large-scale radiology report dataset, 2) query answering based on natural language inference, and 3) interpretable classification based on variational information pursuit.
3.1 Problem Formulation
We consider the problem of classifying a radiology report based on a sequence of interpretable queries. Let $X\in\mathcal{X}$ be a random variable representing a radiology report, where $\mathcal{X}$ is the space of all possible reports, and $Y\in\mathcal{Y}=\{0,1\}$ denote the corresponding binary label. Our goal is to determine $Y$ given $X$ by sequentially querying $X$ with a set $Q$ of predefined, interpretable queries.
Each query $q\in Q$ is a function $q:\mathcal{X}\to\mathcal{A}$, where $\mathcal{A}=\{-1,0,1\}$ represents the possible answers: negative, unknown, and positive. For example, given a radiology report $x$ and a query $q=$Is there a decreased opacity on the right?”, $q(x)=1$ implies there is a decreased opacity on the right, $q(x)=-1$ implies there is not, and $q(x)=0$ implies the query cannot be answered from the given report. We assume the query set $Q$ is interpretable, task-specific, and sufficient [^1] That is, $P(Y\mid x)=P(Y\mid\{x^{\prime}\in\mathcal{X}:q(x)=q(x^{\prime}),q\in Q\})$. See [9]. .
The classification process involves selecting the most informative queries in an iterative manner until sufficient information is obtained to predict $Y$ with high confidence.
3.2 Constructing and Answering Queries from Existing Database
Our final goal is to classify whether a certain disease is present or not in a given CRR $x$. As a prerequisite to applying the IP framework, we first need to define a set of task-relevant and interpretable queries that inform us about the report. In this work, we choose a query that corresponds to a question on whether a particular fact is present in the given report. For example, the fact “The heart shape is normal” can be thought of as a query by rephrasing it as “Did the report mention that the heart shape is normal?” While there are many potential choices for a query set, we argue that curating a sufficiently large set of facts as queries for CRR classification would allow us to build interpretable models with good predictive performance. In what follows, we describe how to construct this query set and how to answer queries for a given radiology report.
Queries $q$. The procedure for obtaining a large set of representative, domain-specific, and interpretable queries is illustrated in Figure 2 Part I. We construct the query set $Q$ by mining facts from a large-scale radiology report dataset using a recently developed fact extraction model [22]. Specifically, we analysis over 220k radiology reports from MIMIC-CXR dataset [15], parse them into over 660k sentences using Natural Language Toolkit [2], and then extract over 590k facts from these sentences. To extract facts from sentences, we first prompt the GPT-3.5 and GPT-4 models to generate facts from each sentence on a subset of sentences, then use these input-output pairs as training data to finetune a sequence-to-sequence model named T5-small [27] for fact extraction. After fact extraction, we encode each fact into a 128-dimension latent space using the fact encoder CXRFE [22]. We apply $k$-means clustering to all 590k extracted facts in the fact embedding space, then select the nearest fact to each cluster center, and conduct deduplication if any two selected facts have a cosine similarity greater than 0.97. After this process, we obtain 520 representative facts, where each of them is a short fact describing the clinical findings, e.g., “moderate size hiatal hernia” or “increased bibasilar opacities”.
Query Answers $q(x)$. Ideally, we hope to directly obtain grounded query answers from the report via annotations from radiologists. However, manual annotation is infeasible due to the large number of reports and queries. Moreover, in contrast to image classification, it often occurs that a query cannot be answered based on the provided radiology report. For example, if the description of the heart is not mentioned in the report, we cannot answer the query “Is the heart size normal?” from the report, although the heart is shown in the image. To obtain the query answers in a scalable and accurate manner, we propose to leverage natural language inference (NLI). Specifically, we pass both the query and the CRR as inputs to an LLM named Flan-T5-large [11] and ask the LLM to predict the query answer as either “yes”, “no” or “unknown”. An example of the LLM prompt is illustrated in Figure 2 Part II. The prompt consists of three parts: instruction, premise, and hypothesis. The instruction serves as guidance to the model and is fixed for every sample, the premise is the radiology report, and the hypothesis is the fact to be verified whether it is true for the given report or not. For every report and query, we set the predicted output of the model as the query answer $q(x)\in\{-1,0,1\}$.
3.3 Classification Based on Variational Information Pursuit
IP. We classify a report using Information Pursuit (IP), an interpretable-by-design framework that sequentially selects queries about the data in order of information gain. Given a sample $x^{\text{obs}}\in\mathcal{X}$, the IP algorithm selects the queries as follows:
$q_{1}=\operatorname*{argmax}_{q\in Q}I(q(X);Y);\quad q_{k+1}=\operatorname*{argmax}_{q\in Q}I(q(X);Y\mid q_{1:k}(x^{\text{obs}})).$ (1)
Here $q_{k+1}\in Q$ refers to the new query selected at step $k+1$ based on the history $q_{1:k}(x^{\text{obs}})$. The algorithm terminates at iteration $L$ when there are no more informative queries, i.e., $\forall q\in Q,I(q(X);Y|q_{1:L}(x^{\text{obs}}))=0$. Furthermore at each iteration $k$, one can also compute the posterior distribution (i.e. the prediction) $P(Y\mid q_{1:k}(x^{\text{obs}}))$. To carry out IP, [9] proposed to learn a VAE to model the joint distribution of all query answers and labels $P(Q(X),Y)$ directly. However, learning a good generative model has shown to be challenging when the size of $Q$ is large, so a variational approach is proposed as an alternative.
V-IP. Rather than modeling the joint distribution of queries and labels to directly compute the most informative queries, Variational Information Pursuit (V-IP) [7, 8, 12] finds a querier $g$ that selects the most informative query $q$ given any random history $S$ of query-answer pairs and a predictor $f$ that models the posterior distribution of $Y$ given $q$ and $S$, by optimizing the following objective:
$\min_{f,g}$ $\quad\mathbb{E}_{X,S}\left[D_{KL}(P(Y\mid X)\ \|\ \hat{P}(Y\mid q(X),S))\right]$ (2)
where $\quad q:=g(S)\in Q$ (3)
$\quad\hat{P}(Y\mid q(X),S):=f(\{q,q(X)\}\cup S).$ (4)
The variational objective aims to minimize the KL-divergence between the true posterior distribution and the posterior distribution given any history and an additional query and its answer. In practice, the classifier and querier are parameterized as neural networks, $f_{\theta}$ and $g_{\eta}$, making them efficient for carrying out the IP algorithm and scalable for evaluating large-scale datasets. The history $S$ is represented using a masked vector with length $|Q|$. Since our task is classification, where $Y$ is a categorical variable, the KL-divergence is empirically equivalent to the cross-entropy loss. Once $f_{\theta}$ and $g_{\eta}$ are trained, the querier $g_{\eta}$ serves as a function that directly outputs the query with the highest mutual information [7], hence IP can be directly carried out by having $q_{k+1}=g(\{q_{1:k},q_{1:k}(x)\})$.
To implement the whole framework of IP-CRR, we first generate a query set which consists of queries from large amounts of radiology reports, then answer each query for each report with an LLM. Finally, classification is conducted by generating a sequence of interpretable query-answers.
4 Experiments
We refer the entire pipeline, from generating queries from the radiology report database, answering queries via natural language inference, and applying V-IP to our task, as IP-CRR. Our proposed framework is modified to handle query answers $q(x)$ obtained from NLI, and our framework naturally handles the issue with unknown query answers. In this section, we will demonstrate its effectiveness in performing a range of tasks in radiology report classification.
4.1 Dataset and Implementation Details
Dataset. To verify the effectiveness of the proposed method, we conduct experiments on a public radiology dataset named MIMIC-CXR [15]. In this dataset, each patient is called “subject”, and each examination is called “study”. Each study corresponds to one radiology report and multi-label ground-truth label indicating whether diseases are present. The dataset is randomly split into 132296, 14914, and 39613 studies separately for training, validation, and testing.
Implementation Details. To train the interpretable model, a querier and a classifier are designed. The architecture of these two components are both five-layer multi-layer perceptions (see code for details). The IP-CRR model terminates with a confidence threshold of 0.85 or a maximum number of queries of 200. The optimizer is AdamW [20] with an initial learning rate of 0.0001 and a cosine annealing decay. The model is trained with batch size 128 and 1000 epochs. The loss function is weighted binary cross-entropy to alleviate the class imbalance. All experiments are implemented in PyTorch 2.5.1 and on an NVIDIA A5000 GPU with a memory of 24.5GB.
Table 1: Predictive performance of IP-CRR and baseline models across six tasks, measured by Average precision (AP, left) and F1 score (right). Higher values indicate better performance. The best results among black-box models are underlined, while the best interpretable model results are bolded.
Methods AP F1
LO CA SD CM PE PN LO CA SD CM PE PN
CXR-BERT (FT-Last) 0.900 0.361 0.969 0.864 0.945 0.449 0.829 0.223 0.912 0.789 0.887 0.449
CXR-BERT (FT-All) 0.984 0.992 0.970 0.964 0.962 0.641 0.987 0.991 0.978 0.982 0.953 0.541
Flan-T5-large 0.527 0.073 0.445 0.380 0.616 0.190 0.663 0.139 0.321 0.543 0.754 0.299
CBM 0.947 0.345 0.934 0.791 0.874 0.432 0.884 0.241 0.853 0.738 0.801 0.431
IP-CRR 0.972 0.578 0.959 0.892 0.925 0.468 0.918 0.350 0.889 0.811 0.860 0.451
(a) Average precision over number of queries asked. Dashed line represents the maximum performance over the number of queries.
(b) Number of queries needed over thresholds on the posterior. Shaded regions indicate the variance across runs.
Figure 3: Quantitative evaluation of IP-CRR as the number of queries increase.
4.2 Quantitative Results and Comparisons
We evaluate the predictive performance of IP-CRR with all queries on six binary prediction tasks: Lung Opacity (LO), Calcification of the Aorta (CA), Support Devices (SD), Cardiomegaly (CM), Pleural Effusion (PE), and Pneumonia (PN). Our baseline methods include 1) CXR-BERT (FT-Last): CXR-BERT [3], with fine-tuning on the last layer; 2) CXR-BERT (FT-All): CXR-BERT [3], with fine-tuning on all layers; 3) Flan-T5-large model [11], without any fine-tuning; and 4) CBM [18]. Notably, 1-3) are in the class of black-box models and 4) is an interpretable-by-design model. The results are listed in Figure 1. Our metric of choice is the Average Precision (AP) and F1 score.
Compared with black-box models, IP-CRR achieves a higher AP and F1 score against CXR-BERT (FT-Last), Flan-T5-large, and CBM for LO, CA, CM, and PN, and achieves comparable performance for SD. However, when compared with CXR-BERT (FT-All), IP-CRR achieves a similar AP for LO, but falls short on other tasks. We believe the difference in performance is largely attributed to CXR-BERT’s black-box nature, its complex architecture, and the fact that it was fine-tuned on BERT. On the contrary, IP-CRR is implemented using a simple MLP with much fewer parameters.
Furthermore, in Figure 3, we evaluate the overall AP as a function of the number of queries. For the task of LO, IP-CRR reaches nearly 0.95 AP with just 30 queries (out of 520 total), and improves to 0.97 with 100 queries – less than 20% of the total number of queries. We find a similar trend for other tasks, which demonstrates that IP-CRR generally selects informative queries in its early iterations. Alternatively, we observe that IP-CRR converges to the maximum performance it can achieve with only a small subset of queries. Compared to baselines in Table 1, IP-CRR achieves a better performance with smaller number of queries than CBMs with all available ones.
Last but not least, in Figure 3, we plot the number of queries needed as the threshold on the posterior changes. The threshold on the posterior approximately measures the confidence the user requires the predictor to reach before terminating the IP-CRR algorithm. We observe that the IP-CRR predictor requires fewer than 150 queries to achieve a confidence of 0.95 for most of tasks. This demonstrates that IP-CRR is able to find informative queries to make confident predictions.
4.3 Qualitative Analysis of Query-answer Chains
Figure 4: Examples of query-answer chains from IP-CRR. Each row in the matrix represents the posterior $P(Y\mid q_{1:k}(x))$ after selecting the query $q_{k}$. The text is $q_{k}$, with red implying the fact is not present $q_{k}(x)=-1$, green implying is present $q_{k}(x)=1$, and black implying it cannot be answered $q_{k}(x)=0$. “N” or “P” means negative or positive prediction of the disease, respectively.
The core mechanics of using IP-CRR for interpretable predictions is the iterative process of selecting informative queries. This allows the user to observe how the posterior evolves as the number of queries increases. In Figure 4, we demonstrate two examples of query-answer chains obtained using IP-CRR. For the example on the left, our method first selects $q_{1}=$Opacification likely reflects atelectasis?”. From the report, the model infers that $q_{1}(x)$ is not present in the report. Then, it selects the query $q_{2}=$Patchy opacity at right lung base?”. Continuing in this fashion, after 9 queries, the model concludes that the patient has LO. The same sequential deduction can be observed in the example on the right, highlighting the interpretability IP-CRR offers when applied to radiology reports.
5 Conclusion
In this work, we propose an interpretable-by-design text classification framework. We obtain a large set of queries by mining existing radiology reports, and leverage NLI for answering these queries. To evaluate the proposed framework, we conducted experiments on a large-scale radiology dataset and on multiple tasks, with both quantitative and qualitative results that illustrate interpretability and competitive performance. As for limitation, our current method relies on ternary query-answers for all queries, and we reserve leveraging stronger Foundation Models in the medical domain for more expressive query-answers as future work.
Acknowledgments
Yuyan, Ryan, and René acknowledge the support from the Penn Engineering Dean’s Fellowship, the startup funds, and the NIH grant R01NS135613. Pablo received support in Chile from the National Agency for Research and Development (ANID) through the Millennium Institute for Intelligent Healthcare Engineering - iHEALTH (ICN2021_004), the National Center for Artificial Intelligence - CENIA (FB210017), and the ANID Scholarship Program / Doctorado Becas Chile / 2019 - 21191569.